A firm has a market capitalization (market value of equity) of $30 Billion and net debt of $4 Billion. Calculate the weight of debt in the firm's weighted average cost of capital (WACC) calculation. 
Answer #1

Weight of debt = Market capitalization of Debt / Total market capitalization of debt and Equity

Weight of debt = $4 billion / ($30 billion + $4 billion)

Weight of debt = $4 billion / ($34 billion)

Weight of debt = 0.1176

or Weight of debt = 11.76%
Answer #2

ANSWER :




C = E + D = 30 + 4 = 34 billion dollars.


Weight of debt in the capital = D/C = 4 / 34 = 0.1176 = 11.76% (ANSWER).

answered by: Tulsiram Garg